OpenStudy (anonymous):

what is the lcm of 12x and 2x then?

OpenStudy (anonymous):

12x

OpenStudy (anonymous):

how do you do it?

OpenStudy (anonymous):

2x is the smaller number i first check if i can multiply a constant (integer) to it to make it 12x 6 * 2x = 12x hence 12x is the LCM

OpenStudy (anonymous):

oh, okay. Thank you. (:

OpenStudy (anonymous):

yup ^_^

OpenStudy (anonymous):

so then how would you do 5(x+y) 15(x+y)

OpenStudy (anonymous):

same thing 5(x + y) is smaller see if there's a constant that can be multiplied to get to 15(x + y) 3 * 5(x + y) = 15(x + y) so it's 15(x + y)

OpenStudy (anonymous):

okay. (:

OpenStudy (anonymous):

cheers ^_^

OpenStudy (anonymous):

Thank you. I don't like math very much.

OpenStudy (anonymous):

lols if you don't like it learn it till you can ace it, then you don't ever need to worry about it good luck~

OpenStudy (anonymous):

Haha, I'll try.